We know that the cos 60 degrees = 1/2. Show that the above is true using a half angle identity.


1
Expert's answer
2021-05-26T15:13:35-0400

"cos(x/2)=\sqrt{(1+cosx)/2}"

for "x=120^o"

"cos(120^{\circ}/2)=\sqrt{(1+cos120^o)/2}=\sqrt{(1-0.5)/2}=\sqrt{0.5/2}=\sqrt{1/4}=1/2"